OpenStudy (anonymous):

10 and 1/3-3 and 1/4*2 and 2/3 A.1 and 1/3 B.1 and 2/3 C.1 and 3/4 D18and8/9

OpenStudy (anonymous):

@q12157

OpenStudy (anonymous):

@LizzyDizzy @livias.random

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (anonymous):

2/3 and 1/3 can be combined to equal 1..

OpenStudy (anonymous):

try typing it into https://mathway.com/ okay hun

OpenStudy (anonymous):

2*1/4=2/4

OpenStudy (anonymous):

which is equal to 1/2

OpenStudy (anonymous):

1+1/2+10-3 is what you are left with

OpenStudy (anonymous):

so it should be 8 and 1/2

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

Are you sure you copied the question correctly?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

10 and 1/3-3 and 1/4*2 and 2/3 this is the correct question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then its none of the answers. are the answer choices copied correctly?

OpenStudy (anonymous):

please check otherwise i cant help you.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/3-3 and 1/4*2 For these two values, can you rewrite them using parentheses?

OpenStudy (anonymous):

it is (10 1/3 - 3 1/4)*2 2/3

OpenStudy (anonymous):

@q12157

OpenStudy (anonymous):

what is equivalent to 8 1/2

OpenStudy (anonymous):

Oh thanks for the parentheses

OpenStudy (anonymous):

Ill do it right now

OpenStudy (anonymous):

combine the stuff in the parentheses to get: (7 1/3-1/4)*2= (7+(1/12))*2=14+1/6 Overall it is 14+(1/6)+(2/3)=14 5/6

OpenStudy (anonymous):

My guess is that you still are missing something...I need more info ><

OpenStudy (anonymous):

the only word missing is evaluate